Im totaly stuck with this one.

( X / 5 ) - ( X + 2 / 3 ) = ( 1 / 15 )

Answer by Nate(3500)

( X / 5 ) - ( X + 2 / 3 ) = ( 1 / 15 )
(( X / 5 ) - X - 2 / 3 ) = ( 1 / 15 ))15
3x - 15x - 10 = 1
-12x = 11
x = -11/12